# Broadening Discovery through Structural Models: Multimodal Combination of Local and Structural Properties for Predicting Chemical Features

### v1

@@ Abstract
In recent years, machine learning has profoundly reshaped the field of chemistry, facilitating significant advancements across various applications, including the prediction of molecular properties and the generation of molecular structures. Language models and graph-based models are extensively utilized within this domain, consistently achieving state-of-the-art results across an array of tasks. However, the prevailing practice of representing chemical compounds in the SMILES format –used by most datasets and many language models – presents notable limitations as a training data format. In contrast, chemical fingerprints offer a more physically informed representation of compounds, thereby enhancing their suitability for model training. This study aims to develop a language model that is specifically trained on fingerprints. Furthermore, we introduce a bimodal architecture that integrates this language model with a graph model. Our proposed methodology synthesizes these approaches, utilizing RoBERTa as the language model and employing Graph Isomorphism Networks (GIN), Graph Convolutional Networks (GCN) and Graphormer as graph models. This integration results in a significant improvement in predictive performance compared to conventional strategies for tasks such as Quantitative Structure-Activity Relationship (QSAR) and the prediction of nuclear magnetic resonance (NMR) spectra, among others.
@@ 1 Introduction
The integration of machine learning (ML) has emerged as a transformative force in the natural sciences, particularly in the discipline of chemistry [Chithrananda et al., 2020, Hu et al., 2016, Wang et al., 2022]. This integration encompasses various tasks, ranging from the regression of molecular properties, exemplified by quantitative structure-activity relationship (QSAR) models [Wu et al., 2021, Rácz et al., 2021], to complex challenges, such as predicting nuclear magnetic resonance (NMR) spectra from structure of chemical compound [Yao et al., 2023]. As an ever-evolving discipline, the latest advancements in machine learning are gradually being adapted for applications in chemistry, albeit with some delay. Molecular representations are fundamental to the application of machine learning in chemistry, and three primary types are typically employed: graph-based [Reiser et al., 2022], string-based [Heller et al., 2015, Weininger, 1988, Krenn et al., 2020], and vector representations [Rogers and Hahn, 2010, Durant et al., 2002].
@@
Graph-based representations conceptualize chemical compounds as molecular graphs, effectively capturing their structural properties [David et al., 2020, Kwon et al., 2020]. This format naturally aligns with graph neural networks, which have been successfully applied to numerous chemical problems, demonstrating their efficacy in molecular analysis. String representations, particularly Simplified Molecular Input Line Entry System (SMILES) [Weininger, 1988], are widely regarded as a standard method for the linear representation of molecular structures. SMILES is typically used for storing compounds in databases and, despite its limitations, effectively represents the structure of a molecule [Chithrananda et al., 2020, Ahmad et al., 2022, Wang et al., 2019, Shamshad et al., 2023, Cong et al., 2024]. However, it presents notable shortcomings. Initially designed for efficient storage and representation of molecular data, SMILES lacks comprehensive information regarding the physical and chemical properties of compounds. As machine learning progresses, the inadequacies of SMILES in facilitating in-depth semantic analysis have become increasingly evident. Recent methods have sought to combine graph models with SMILES-based natural language processing (NLP) transformers [Zhu et al., 2023], thereby integrating the strengths of both methodologies.
The other type of representation comprises vector representations, notably fingerprints, which were developed for substructure identification and similarity searching. Numerous studies, such as those presented by [Sabando et al., 2022] and [Wang et al., 2020], have highlighted the effectiveness of fingerprints as molecular representations in machine learning tasks, demonstrating promising results even with conventional algorithms and simple models. However, the potential of fingerprints (including ECFP) in the context of modern language models remains largely underexploited.
We believe this approach is particularly advantageous due to the transformer’s inherent versatility and flexibility, allowing it to address a variety of physicochemical tasks that differ significantly in nature. For example, fingerprints-based architecture could potentially handle tasks such as predicting molecular property values [Wu et al., 2021, Rácz et al., 2021, Sabando et al., 2022], classifying compounds based on their biological activity, generating novel molecular structures, exploring molecular interactions [Hu et al., 2016] and complex challenges, such as predicting nuclear magnetic resonance (NMR) spectra from molecular structures [Kwon et al., 2020] or solving the co-crystallisation problem [Wang et al., 2020].
Given the diverse challenges presented by these tasks, we seek to create a unified framework that requires only minor modifications for each specific application. By leveraging the strengths of transformer models, we anticipate that our architecture will enable efficient and robust performance across multiple domains within the field of cheminformatics.
The questions we want to tackle with our research are the following:
  1. To what extent can a language transformer model based on fingerprints enhance performance compared to conventional approaches?
  2. How much results can be improved by combining such language transformer with graph model?
  3. Furthermore, can transformers trained on fingerprint representations improve the quality of multitasking embeddings, thereby providing more robust and nuanced representations that capture the complexities across diverse tasks?
@@ 2 Related Work
2.1 Extended-Connectivity Fingerprints
Extended-Connectivity fingerprints [Rogers and Hahn, 2010], are so-called circular fingerprints that assign a two-dimensional hash array to each molecule. Each element of such an array is a hash corresponding to one atom. It encrypts a fixed set of physical and chemical properties of this atom, such as charge, as well as information about its neighbours.
As applied to our task, there are three significant particularities of ECFP. Firstly, the fingerprint of a single molecule consists of an array of hashes (i.e. in NLP terms, we can think of the array as a sentence and each individual hash as a word). Secondly, each hash is constructed based on a set of physical properties. Thus, each array element is based on physical and chemical data. And lastly, there is a so-called diameter, which shows neighbouring atoms in one iteration. That being said, we cover only those atoms, that are within the diameter’s reach. This sensible of surrounding environment representation can be very useful for such tasks as molecular NMR spectroscopy, where chemical environment plays crucial role in spectrum definition.
Over the last few years, a number of methods [Wang et al., 2020, Sturm et al., 2018] have pointed out the effectiveness of using ECFP as features for training quite simple models (such as SVM) to solve various chemical problems.
Several approaches have employed ECFP as a representation for training data in natural language processing (NLP) algorithms; however, these methods predominantly rely on relatively conventional machine learning techniques. One notable example is Mol2Vec [Jaeger et al., 2018], which implements the Word2Vec algorithm utilizing ECFP data representation.
2.2 SMILES-based NLP models
@@
Transformers [Vaswani, 2017] were initially introduced to facilitate the generation of vector representations for natural language processing tasks. Since their inception, they have found widespread application across a variety of domains, including speech recognition, medicine, and neuroscience [Shamshad et al., 2023, Cong et al., 2024].
There have been several efforts to adapt transformers for chemical applications, exemplified by models such as SmilesBERT [Wang et al., 2019], ChemBERTa [Chithrananda et al., 2020], and ChemBERTa-2 [Ahmad et al., 2022], and all of them were trained on SMILES.
@@
Many of these models have been trained on substantial datasets, including ZINC [Irwin et al., 2012] and PubChem [Kim et al., 2023], demonstrating commendable performance in classification and regression tasks across various established chemical benchmarks.
2.3 Graph models
@@
Graph neural networks (GNNs) have been effectively utilized to address a variety of challenges within the field of chemistry [David et al., 2020, Kwon et al., 2020]. Many GNNs are highly specialized for specific tasks and are not inherently designed for generating vector representations of chemical compounds.
@@
In the graph component of our model, we advocate for an approach that synthesizes these concepts and leverages state-of-the-art models. Specifically, we implement a mechanism to mask atom features and edge features in the case of Graphormer [Ying et al., 2021]. The model is trained not only to predict these masked features but also to align the embeddings of two augmented versions of the same molecule. This approach represents a modification of contrastive learning, a technique that remains underutilized in the chemistry domain.
Moreover, [Zhu et al., 2023] introduced a bimodal architecture incorporating a BERT-based language model (LM) trained on SMILES alongside a GNN as the graphical representation model. In contrast, we propose a distinct language model that is trained on fingerprints, thus providing a more physically informed perspective and an advanced graph model. Additionally, our approach includes notable differences in the final projection and the processing of embeddings derived from both the language and graph models.
3 Our contributions
In this paper, we propose a novel methodology [^1] Our code for all experiments is accessible on https://anonymous.4open.science/r/Transformers-for-Molecules-D10E. that integrates graph-based representations with language models based on fingerprints, effectively addressing these limitations. Our approach encompasses four distinct architectures.
RoBERTa with ECFP.
The model implemented in this study is based on the RoBERTa framework and employs ECFP as a baseline for our investigations. As previously discussed, ECFP possesses a fixed radius, rendering it particularly effective for capturing the local properties of substructures or individual atoms within chemical compounds. Both ECFPs and language models effectively leverage their respective training data to extract local properties, showcasing their strength in identifying meaningful features within a localized context. This parallel emphasizes their utility, as noted by Iman Mirzadeh et al. [Mirzadeh et al., 2024], where language models similarly prioritize local information, while have some problems with reasoning. The incorporation of a specified radius in ECFPs allows for the adjustment of substructure sizes, which is especially pertinent for a range of chemical tasks, including the prediction of NMR spectra and the analysis of reaction centers.
LM and Graph model.
In contrast, a graph-based model offers a comprehensive representation of the molecular structure, which proves advantageous for analyzing extensive connections characterized by numerous substructures. Such scenarios are frequently encountered in the field of biochemistry, particularly when addressing pharmacological compounds or polymer-related challenges. In this study, we implement atom masking and graph augmentation techniques that more accurately reflect physical properties, thereby enhancing the model’s capacity to understand molecular characteristics and ensuring a robust representation of the chemical structure. Furthermore, in the context of Graphormer, we employ an advanced architecture designed to capture intricate relational patterns within molecular graphs.
We introduce two bimodal architectures that integrate the RoBERTa model as the linguistic component with graph convolutional networks (GCNs) [Kipf and Welling, 2016] and Graph Isomorphism Networks (GINs) [Xu et al., 2018], respectively. These architectures utilize contrastive learning to improve feature extraction. While both models demonstrate faster training times, they may exhibit limitations in their performance on complex tasks when compared to a more sophisticated architecture that combines RoBERTa with Graphormer [Ying et al., 2021], a graph transformer specifically developed to capture intricate relational patterns.
In each of the graph models – GCN, GIN, and Graphormer – we incorporate a mechanism that synthesizes existing concepts through the masking of atom and edge features. Each model is trained not only to predict these masked features but also to align the embeddings of two augmented representations of the same molecule. This approach constitutes a modification of contrastive learning, a technique that remains underutilized within the chemistry domain.
By leveraging both the structural attributes of molecules and the semantic richness of fingerprints, our innovative bimodal architecture significantly enhances predictive performance and facilitates applicability across a diverse array of physicochemical tasks.
4 Method
@@ 4.1 Architecture Overview
The proposed model comprises three primary components, as illustrated in Figure 1: the graph model, the language model, and the projection blocks. The language model is designed to accept ECFP connectives as input, whereas the graph model processes molecular graphs.
The function of the projection blocks is to transform the embeddings generated by the graph and language models from their respective latent spaces into a unified third latent space.
@@ 4.2 Language model
Tokenizer.
At first, we attempt to use hash values from the ECFP format as the direct input of the language model. This idea is not prosperous because the range of the hash function’s outputs (approximately from $-2^{32}$ to $2^{32}$) is too wide to utilize them as tokens for the model’s input. In this case, we decide to include a tokenizer in the processing. This step allows us to narrow down the range of possible model vocabulary values. As we work with an array of integers interpreted as text, we cannot afford to use a normal tokenizer, which creates tokens from text. In this regard, we choose the Byte-Pair Encoding Tokenizer, as shown in Figure 2, which allows the production of tokens from raw bytes. We train this tokenizer on the largest dataset we have – PubChem [Kim et al., 2023], containing 10 million molecules. This pipeline modification decreases the $vocab$-$size$ of the model to not more than $30,522$.
RoBERTa training.
We utilize the RoBERTa architecture [Liu, 2019], which has been trained on ECFPs derived from the PubChem [Kim et al., 2023] and ZINC [Irwin and Shoichet, 2005] datasets, as our language model. Within this framework, the encoding of an individual atom in ECFP is interpreted as a "word," while the encoding of an entire molecule is considered analogous to "text." During the training process, the ECFP undergoes standard procedures including the masking of 15% of tokens (representing atom hashes), with the model subsequently predicting the probabilities of these masked tokens. The output embedding is derived from the CLS token located in the penultimate layer of the model.
Figure 2: Example of tokenization process. Toketns "0" and "2" correspond to BOS (begin of sequence) and EOS (end of sequence), respectively.

### v2

# Thinking like a CHEMIST: Combined Heterogeneous Embedding Model Integrating Structure and Tokens

@@ Abstract
Representing molecular structures effectively in chemistry remains a challenging task. Language models and graph-based models are extensively utilized within this domain, consistently achieving state-of-the-art results across an array of tasks. However, the prevailing practice of representing chemical compounds in the SMILES format – used by most data sets and many language models – presents notable limitations as a training data format. In this study, we present a novel approach that decomposes molecules into substructures and computes descriptor-based representations for these fragments, providing more detailed and chemically relevant input for model training. We use this substructure and descriptor data as input for language model and also propose a bimodal architecture that integrates this language model with graph-based models. As LM we use RoBERTa, Graph Isomorphism Networks (GIN), Graph Convolutional Networks (GCN) and Graphormer as graph ones. Our framework shows notable improvements over traditional methods in various tasks such as Quantitative Structure-Activity Relationship (QSAR) prediction.
@@ 1 Introduction
The integration of machine learning (ML) has emerged as a transformative force in the natural sciences, particularly in the discipline of chemistry [Chithrananda et al., 2020, Hu et al., 2016, Wang et al., 2022]. This integration encompasses various tasks, ranging from the regression of molecular properties, exemplified by quantitative structure-activity relationship (QSAR) models [Wu et al., 2021, Rácz et al., 2021], to complex challenges, such as predicting nuclear magnetic resonance (NMR) spectra from the structure of chemical compounds [Yao et al., 2023]. As an ever-evolving discipline, the latest advancements in machine learning are gradually being adapted for applications in chemistry, albeit with some delay. Molecular representations are fundamental to the application of machine learning in chemistry, and three primary types are typically employed: graph-based [Reiser et al., 2022], string-based [Heller et al., 2015, Weininger, 1988, Krenn et al., 2020], and vector representations [Rogers and Hahn, 2010, Durant et al., 2002].
@@
Graph-based representations conceptualize chemical compounds as molecular graphs, effectively capturing their structural properties [David et al., 2020, Kwon et al., 2020]. This format naturally aligns with graph neural networks, which have been successfully applied to numerous chemical problems, demonstrating their efficacy in molecular analysis. String representations, particularly Simplified Molecular Input Line Entry System (SMILES) [Weininger, 1988], are widely regarded as a standard method for the linear representation of molecular structures. SMILES is typically used for storing compounds in databases and, despite its limitations, effectively represents the structure of a molecule [Chithrananda et al., 2020, Ahmad et al., 2022, Wang et al., 2019, Shamshad et al., 2023, Cong et al., 2024] and serves as a basic data representation for language models [Chithrananda et al., 2020, Ahmad et al., 2022, Ross et al., 2022]. However, it presents notable shortcomings [Ganeeva et al., 2024]. Initially designed for efficient storage and representation of molecular data, SMILES lacks comprehensive information regarding the physical and chemical properties of compounds. Two main approaches exist to overcome these difficulties. The first combines graph models with SMILES-based natural language processing (NLP) transformers [Zhu et al., 2023], integrating the strengths of both methodologies.
The second is based on changing the representations of molecules used in training. In cheminformatics, molecules are commonly represented as vectors in a high-dimensional space, where each vector encodes essential molecular features to enable effective computational analysis. Molecular descriptors [Sahoo et al., 2016, Bhatia et al., 2023, Darlami and Sharma, 2024] – numerical values summarizing properties such as size, shape, and electronic characteristics – are central to this representation, providing critical information for predicting molecular behavior and interactions. Traditional global descriptors capture information from the entire molecule but often fail to reflect important structural variations within specific subregions. To address this limitation, we propose focusing on descriptors derived from molecular substructures, offering a more detailed and accurate representation that better captures the complexity of molecular behavior. There are existing attempts to use quantum-chemical descriptors as a basis for transformer-based models [Zhou et al., 2023].
Thus, there is a legitimate idea of combining these two approaches: adapting descriptors as a data format for language models and applying bimodal architectures.
Contributions.
We propose an approach that involves decomposing molecules into chemically meaningful substructures and calculating descriptors for each segment. These sequences of substructure descriptors serve as input to the RoBERTa model [Liu, 2019], enabling it to learn and capture the underlying physicochemical relationships within the molecule. This helps in improving performance in tasks such as predicting molecular properties [Wu et al., 2021, Rácz et al., 2021, Sabando et al., 2022], classifying biological activity, generating novel molecules, and studying molecular interactions [Hu et al., 2016].
The graph-based model provides a detailed representation of molecular structures, making it especially effective for analyzing large compounds with multiple substructures. By explicitly capturing the connectivity and spatial relationships among substructures, it overcomes the limitations of language models that lack detailed molecular organization. To improve structural representation and property prediction, we use atom masking and graph augmentation with Graphormer, an advanced architecture designed to capture complex relational patterns in molecular graphs.
We propose two bimodal architectures combining RoBERTa with a graph convolutional network (GCN) [Kipf and Welling, 2016] and a Graph Isomorphism Network (GIN) [Xu et al., 2018], using contrastive learning to enhance feature extraction. Although these models train faster, they may underperform on complex tasks compared to a more advanced system integrating RoBERTa with Graphormer [Ying et al., 2021], which specializes in modeling intricate relations. In all graph models, we apply masking of atom and edge features during training to predict masked elements and align embeddings of augmented molecular views. This contrastive learning approach is underexplored in cheminformatics and shows promise for advancing molecular modeling. Code is available F.
@@ 2 Related Work
Molecular descriptors. The relationships between molecular properties have been extensively studied and form a foundation for property prediction in chemoinformatics [Hansch et al., 1962, 1963, Hansch and Fujita, 1964]. Molecular descriptors, which quantitatively represent molecular structure and characteristics, encompass various types such as substructural descriptors (e.g., MACCS keys), topological descriptors derived from 2D molecular graphs, and geometric descriptors capturing 3D molecular shape and spatial configuration. These descriptors serve as key inputs for QSAR/QSPR models that aim to predict molecular properties from structural information (appendix A). While conventional models typically rely on whole-molecule descriptors, they often overlook the nuanced influence of individual substructures, particularly in larger, multifunctional molecules [Estrada, 2008]. To address this limitation, recent strategies focus on fragment-based descriptors, enabling more localized and interpretable predictions.
When applied as representations for language models, molecular descriptors offer promising advantages. Certain descriptors encode positional information reminiscent of natural language sequences while simultaneously reflecting important physicochemical properties. This dual capability allows language models to capture both statistical patterns and chemically meaningful relationships, improving their ability to generate insightful and reliable predictions.
BRICS fragmentation. The Breaking of Retrosynthetically Interesting Chemical Substructures (BRICS, appendix B) method [Degen et al., 2008] offers a principled, rule-based approach to fragmenting molecules into chemically meaningful components by selectively cleaving bonds commonly involved in synthesis, guided by 16 well-defined rules. This ensures that the resulting fragments correspond to synthetically relevant building blocks while retaining information about their potential attachment points, facilitating their recombination in silico. Widely adopted in drug discovery and computational chemistry, BRICS mirrors the logic of retrosynthetic analysis by breaking down complex molecules into simpler precursors.
In our work, we utilize BRICS to preprocess data by partitioning molecules into fragments and computing descriptors for each. This fragment-centric strategy shifts focus from the whole molecule to its local chemical environments, improving the model’s ability to predict properties influenced by specific substructures and enabling a more interpretable analysis of structure–property relationships.
@@
SMILES-based NLP models. Transformers [Vaswani, 2017] were initially introduced to facilitate the generation of vector representations for natural language processing tasks. Since their inception, they have found widespread application across various domains, including speech recognition, medicine, and neuroscience [Shamshad et al., 2023, Cong et al., 2024].
Figure 1: An example of splitting molecule to substructures and creating the additional substructure (in case of serotonin, its the parent molecule due to it was splitted to only two BRICS blocks).
@@
There have been several efforts to adapt transformers for chemical applications, exemplified by models such as SmilesBERT [Wang et al., 2019], ChemBERTa [Chithrananda et al., 2020], and ChemBERTa-2 [Ahmad et al., 2022], and almost all of them were trained on SMILES. Many of these models have been trained on substantial datasets, including ZINC [Irwin et al., 2012] and PubChem [Kim et al., 2023], demonstrating commendable performance in classification and regression tasks across various established chemical benchmarks.
@@
Graph models. Graph neural networks (GNNs) have effectively addressed a variety of challenges within the field of chemistry [David et al., 2020, Kwon et al., 2020]. Many GNNs are highly specialized for specific tasks and are not inherently designed to generate vector representations of chemical compounds.
@@
In the graph component of our model, we advocate for an approach that synthesizes these concepts and leverages state-of-the-art models. Specifically, we implement a mechanism to mask atom features and edge features in the case of Graphormer [Ying et al., 2021]. The model is trained not only to predict these masked features but also to align the embeddings of two augmented versions of the same molecule. This approach represents a modification of contrastive learning, a technique that remains underutilized in the chemistry domain. Moreover, [Zhu et al., 2023] introduced a bimodal architecture incorporating a BERT-based language model (LM) trained on SMILES alongside a GNN as the graphical representation model. In contrast, we propose a distinct language model that is trained on fingerprints, thus providing a more physically informed perspective and an advanced graph model. Additionally, our approach includes notable differences in the final projection and the processing of embeddings derived from both the language and graph models.
3 Data Preprocessing
We propose a data preparation methodology for language models that focuses on representing local substructural properties through physicochemical descriptors of molecules. The process divides into several key stages. First, molecules represented by SMILES are partitioned into substructures using the BRICS algorithm. Then, additional substructures are generated: for every bond removed during this fragmentation, a new substructure is created that comprises the two BRICS-derived substructures originally connected by that bond (see Fig. 1). To maintain chemical completeness, all substructures are augmented with hydrogen atoms according to valence rules, compensating for broken bonds in the parent molecule. This design, somewhat analogous to circular fingerprints such as ECFP [Rogers and Hahn, 2010] or Morgan fingerprints [Morgan, 1965], enables the model to capture information not only about individual substructures but also about the connections between them.
Figure 2: Example of tokenization process. Toketns "0" and "2" correspond to BOS (begin of sequence) and EOS (end of sequence), respectively. The ’!’ and ’\$’ kept non-tokenizenised for clarity.
Subsequently, a set of structural, topological, and physicochemical descriptors (described in the appendix A) is calculated for each substructure via RDKit. These descriptors are organized into an ordered array, which is crucial for incorporating positional information and the usage of positional encoding in RoBERTa). The selected descriptors capture various aspects, including the topological structure of the molecule, the number of atoms in each substructure, as well as certain quantum and physicochemical properties (see the first two arrows in Fig. 5). As a result, each molecule is transformed first into a collection of relevant substructures and is represented as a two-dimensional array, where each row corresponds to the descriptor vector of an individual substructure (see the next section).
Discussion. Our approach relies on partitioning molecules into chemically meaningful substructures using the BRICS method, rather than applying conventional BPE-like tokenization, which often emphasizes frequently occurring but not necessarily chemically relevant fragments. Importantly, the BRICS fragmentation is guided by a set of rules specifically developed for retrosynthetic analysis, mirroring the way chemists cognitively organize molecular structures into building blocks when designing new compounds or predicting their properties. Since chemistry as a scientific discipline is fundamentally rooted in human reasoning and intuition, enabling the model to understand molecules through these chemically intuitive substructures allows it to capture the underlying chemical principles more effectively. By aligning the model’s perspective with the way chemists think, rather than relying solely on statistical patterns, we improve its ability to grasp the chemical laws and concepts established through decades of human expertise, thereby enhancing both interpretability and practical relevance.
It is also important to recognize that many molecular properties cannot be accurately represented as a simple linear combination of the properties of individual substructure descriptors. This challenge is analogous to the problem of constructing a meaningful sentence embedding from the embeddings of its constituent words. In this context, substructures can be viewed as words, while the entire molecule corresponds to a sentence, with the number of substructures varying between molecules. The RoBERTa architecture proves to be particularly well-suited for addressing this type of variable-length, context-dependent representation, making it an optimal choice for modeling molecular properties in this framework.
4 Architecture and training details
@@ 4.1 Architecture Overview
The proposed model comprises three primary components, as illustrated in Figure 3: the graph model, the language model, and the projection blocks. The language model is designed to accept two-dimensional arrays of substructure descriptors as input, whereas the graph model processes molecular graphs. The function of the projection blocks is to transform the embeddings generated by the graph and language models from their respective latent spaces into a unified third latent space. The baseline of our approach is the language model itself (without graph parts), and as shown in Section 5, it outperforms other modern frameworks in most benchmarks.
@@ 4.2 Language model
Tokenizer. Our input to the language model consists of two-dimensional arrays of descriptors, where each array represents a single molecule and each subarray corresponds to a specific substructure within that molecule. The tokenization process applied to these arrays is relatively straightforward, as the nature of the chosen descriptors eliminates the need for conventional statistical tokenization methods, such as Byte Pair Encoding (BPE), which aim to adjust the dictionary size.
Instead, the tokenization involves four key steps. First, within each subarray, the descriptor at position $i$ is adjusted by adding the sum of the maximum possible descriptor values from the preceding positions $[0,i-1]$. This operation effectively tokenizes the descriptor values in a position-aware manner (see the last two arrows on Fig. 2). Next, a special separator token, denoted here as ’!’, is inserted between descriptor subarrays to clearly signal the boundary between different substructures. To further highlight the chemical significance of the fragmentation, an additional token ’\$’ is inserted to separate traditional BRICS-derived substructures from the so-called additional substructures (as detailed in the Data Preprocessing). Finally, the resulting two-dimensional array is flattened into a one-dimensional sequence, with tokens indicating the start and end of the sequence appended at the respective positions.
RoBERTa training. We utilize the RoBERTa architecture [Liu, 2019], which has been trained on descriptors for molecules derived from the PubChem [Kim et al., 2023] dataset, as our language model. Within this framework, the encoding of an individual descriptor is interpreted as a "word", while the substructure in a molecule is interpreted as a "sentence" and lastly the encoding of an entire molecule is considered analogous to "text". During the training process, the 1-dimensional array obtained after the tokenization process undergoes standard procedures, which was given after tokenization process, undergoes standard procedures including the masking of 15% of tokens (representing descriptors), with the model subsequently predicting the probabilities of these masked tokens. The output embedding is derived from the CLS token located in the penultimate layer of the model.